No. 4
Sprint int . 8:01 PM С 46% An experiment to compare the spreading rates of five different brands of yellow interior latex paint available in a particular area used 4 gallons (4) of each paint. The sample average spreading rates (ft2/gal) for the five brands were x1,-462.0, x2. -502.8, X3. 427.5, x4.-469.3, and xs.-532.1. The computed value of F was found to be significant at level a 0.05. With MSE 410.8, use Tukeys procedure to investigate significant differences between brands. (Round your answer to two decimal places.) Which means differ significantly from one another? (Select all that apply.) O x1, and x2. X1. and X3. O X1 and X4 X1 and x5. C X2, and X3. 0-2, and-4. X2, and X5 x3, and X4. O x3, and xs. x4, and x5. . O There are no significant differences. You may need to use the appropriate table in the Appendix of Tables to answer this question. Show My Work (optiona) o Consider the following summary data on the modulus of elasticity (x 106 psi) for lumber of three different grades. Grade 163 29 2 7156 0.23 3 7 142 25 Compute a 95% t CI for e«uan *au-n (Round your answers to three decimal places.) x 106 psi
Answer #1

4:

Here we have 5 different paints so k=5

Total number of data values is n = 4*5 = 20

Degree of freedom of error: df = 20 -5 = 15

Critical value for a-0.05, df=15 and k=5 is

= 4.37

So Tukey's HSD will be

HSD=q\sqrt{\frac{MSE}{2}\left ( \frac{1}{n_{i}}+\frac{1}{n_{j}} \right )}=4.37\sqrt{\frac{410.8}{2}\left ( \frac{1}{4}+\frac{1}{4} \right )}=44.29

Following table shows the absolute difference between all possible pairs of mean:

Samples Samples mean Absolute Difference
1 2 462 502.8 40.8
1 3 462 427.5 34.5
1 4 462 469.3 7.3
1 5 462 532.1 70.1
2 3 502.8 427.5 75.3
2 4 502.8 469.3 33.5
2 5 502.8 532.1 29.3
3 4 427.5 469.3 41.8
3 5 427.5 532.1 104.6
4 5 469.3 532.1 62.8

Following table shows the differences that are greater than Tukey's HSD:

Samples Samples mean Absolute Difference
1 5 462 532.1 70.1
2 3 502.8 427.5 75.3
3 5 427.5 532.1 104.6
4 5 469.3 532.1 62.8

Correct options are:

X1 and X5

X2 and X3

X3 and X5

X4 and X5